Shift report 20011222 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Dec 22 2001 - 16:00:58 EST

  • Next message: Apache: "Shift report 20011223 16:00-24:00"

    ------------------------------------------------------------
    Date: Sat Dec 22 16:00:58 2001, Shift: 08:00-16:00
    Supervisor: fv, with: fv
    Log book#11 pages 74 - 78
    
    Shift Summary:
    ------------------------------------------------------------
    1 .Machine mainly in study mode - so much for physics run.
    
    2. misc clean up of trigger and checkout. see details below
    
    3. The pii4 has no virtual desktop (icons disapeear)
       scaledoan problesm for some trigfger channels - moved around.
    
    Plan: move BFS to align with FFS at 8 deg.
          collect data with Inel+ MRS and FS spec trigger
          at 1/8 FS field and 2Kg MRS
          Aim ~ 50k FS trigger, and similar 2* for MRS.
          before checking response.
    
    8.40: Rhic polarization measurements.
          (Beam should be dumped afterwards).
    
    
    
    Trigger 1 ( Beam-Beam Left+Right all tubes >=1 hit in coinc)
    
    Trigger (Inel )
    
    
    9.00  When they do polarization measurements the inel background 
    rate goes up by ~30% (3000/sec -> 4200/sec).MCR calls ahead of this, Beware that data runs MUST be stopped when Pol measurements are done. 
    It takes ABout 1/2 hour and is done and start and beginning of each store. 
    It is not necessary to turn off/down the anode voltages.
    
    9.10 The printer demon. Seem not to like certain monitor output printerfiles
         and dies. I had to restart lpd. This can only be done by super-user.
         Is the possible reason that the .ps files supermon produces all have the same files??
    
    In addition it gives an error (Betty notice)
    ./lpd start
    Starting lpd: Warning - cannot stat 'if' filter '/var/spool/lpd/psct510a/filter' - No such file or directoryWarning - cannot stat 'if' filter '/var/spool/lpd/pscp510a/filter' - No such file or directory
    
    
    Another comp note:
    On pii4 the virtual desktop bar has disapeared such the iconizes a window 
    mean you cannot get at it again !!!
    Does anyone know how to fix this ?
    
    9.56
    	Ted does not display t1, T2 tracks any longer even when tpc3d shows
    very reasonable clusters, and tpm1 tpm2 does. My guess is it has to do with
    hidden options /requieremnents (like beam beam,) in the tracking package.
    This after After spending (wasting) some time attempting to see why Ted did not display any T1,T2 information though triggering + anodes + etc looks good 
    
    ~10.00 Beam was finally dumped. The polarization measurements took a hole hour
    not so good.
    
    11.20 Machines studies and access at 2.00 (Phenix planned). Beam possibly at
    4.00 . So this is physics running ?? - studies as well as low intensities?
    
    
    11.35 Note Hiro made a new ntuple macro for the pp run it is called
          SummaryPlotsPP.C; runs as usual.
    
    11.45 Modified SummaryPlotsPP.C to only show Tpc sequences 
          for trigger 3 and 6.
    
    12.00 By looking at the ntuple plots from last night it is clear
     the TPC anode were in standby, not on . This explains the lack of 
    tracks correlated to the trigger 3/6.
    
    13.00 Tracking down the 'scale-down problem for trigger 2/4 for the inel 
          counter. Found that the inputwidth to the enable gate 
          (delivered by the top 821 ch #2) was narrower  than for the other inputs.
          This though turns out not to be the problem. It must have to do with the
          nim/ecl-scaledown and back to Nim.
          Setting Inel to channel 5 inttead for long term stability test.
    
    
         Thus from run 6113 and on.
    
         Trigger 1 -- Beam/Beam L.R
         Trigger 3 -- MRS-F.Mrs.Back and Inel
         Trigger 5 -- Inel
         Trigger 6 -- FS D1Tr and H1
    
    13.15 The first majority output of the first Malu (left) has appeared 
          intermittent for some time. Terminated the output and diasychained the
          small & Big tube to same linear amp.
          For the Au-Au the two were independently to seperate fan in/out.
    
    14:14 Clean up of trigger cable delays for MRS and FS trigger.
          All times kept. 
          Still needs fine adjustments.
    
    15:15 1 hours controlled access.
    
    
    -- Beam hours: 8-9 when announchment of pol measurements were made.
    
    
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Dec 22 2001 - 16:01:44 EST